HETAV PATEL

Grade 11,

What is Claisen condensation? Give it’s mechanism and some examples.

What is Claisen condensation? Give it’s mechanism and some examples.

Grade:12th pass

1 Answers

Jayant Kishore
8 years ago
IN CLAISEN CONDENSATION ESTERS REACT WITH R- O- Na FOLLOWED BY H+ TO FORM β KETO ESTERS.
Formation of enolate ion ie removal of α H
C2H5-(C=O)-O-CH3 + CH3 ONa → CH3CH --(C=O)-CH3  
Attack of enolate ion on othe molecule 
CH3CH --(C=O)-CH3  +C2H5-(C=O)-O-CH3     → CH3-(CH-COCH3)-CO-C2H
CHO Na again removes acidic H of active methylene group 
CH3-(CH-COCH3)-CO-C2H+ CHO Na → CH3-(CH-COC -H2)-CO-C2H5
CH3-(CH-COC -H2)-CO-C2H5 + H+ → CH3-(CH-COCH3)-CO-C2H β KETO ESTER
I know mechanism may not be so clear , (I’m not finding the option to attach an image i’ll do it as soon as i ‘ll find it) . When enolate ion attacks on carbonyl gp of other molecule the ester part ie -O – CH3 part of the molecule leaves as it is a good leaving group . The reaction can also be intramolecular if one more carbony group is present in the same molecule. Intramolecular claisen condensation  is prefferd when 5-6 member ring forms.